OpenStudy (anonymous):

@ganeshie8 @theone5526 @Loveiskey18 At the beginning of this month, the balance of Clayton's checking account was $337.55. So far this month, he has received a paycheck via direct deposit of $687.22, been charged a monthly service fee from his bank of $15.00, used a debit card linked to his account to make a purchase of $62.11, written a check for $88.14 that has already been deposited, and deposited a check written to him for $175.00. What is the current balance of Clayton's checking account? A. $1763.18 B. $1888.25 C. $1619.30 D. $1034.52 @satellite73 @Mertsj @Owlcoffee @sierraabel

OpenStudy (owlcoffee):

First, read the problem and interpretate it mathematically.

OpenStudy (anonymous):

i did im confused

OpenStudy (owlcoffee):

okay so, we start with 337.55, how much should we add?

OpenStudy (anonymous):

the answer is 1034.52 so you add 337.55+682.22=1019.77-15.00=1004.77-62.11=957.66-88.14=869.52+175=1034.52

OpenStudy (anonymous):

can you give me a medal

OpenStudy (anonymous):

thank you:)